OpenStudy (anonymous):

how do i find an equation for the tangent to the graph y = arcsecx at x=2

OpenStudy (anonymous):

First find dy/dx = ???

OpenStudy (anonymous):

Are you sure it's x=2?

OpenStudy (anonymous):

yes thats what it says..

OpenStudy (anonymous):

Oh... okay I was reading that wrong... so can you find dy/dx = ???

OpenStudy (anonymous):

Do you need help finding the derivative?

OpenStudy (anonymous):

\[\frac{ 1 }{ \left| x \right| \sqrt{x^2-1}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So you can find the slope using the derivative... Find the slope of the line at x=2

OpenStudy (anonymous):

slope = \(\large m = \frac{ 1 }{ \left| 2 \right| \sqrt{(2)^2-1}}=\frac{1}{2\sqrt3} \)

OpenStudy (anonymous):

Now all you need is the y-coordinate when x=2 So you'll then have a point and a slope to write the equation of the tangent line in point-slope form

OpenStudy (anonymous):

okay hang onn

OpenStudy (anonymous):

well its the same thing lol